OpenStudy (anonymous):

Prob1c, RemainingBalance = (RemainingBalance*(1 + MonthlyRate)) - Payment Gives me 26956.44 for a monthly payment, but test case in handout shows 29643.05

OpenStudy (turingtest):

please show your whole code

OpenStudy (anonymous):

## a credit card program that calculates the minimum fixed monthly payment needed in order to pay off a credit card balance within 12 months. ## this more complicated version uses a bisection search method and upper and lower bounds InitialBalance = float(raw_input('What is the outstanding balance?')) apr = float(raw_input('What is the annual interest rate?')) MonthlyRate = apr/12.0 PaymentLow = InitialBalance/12.0 PaymentHigh = (InitialBalance * (1 + MonthlyRate)**12.0)/12.0 Payment = (PaymentLow + PaymentHigh)/2.0 RemainingBalance = InitialBalance while ((RemainingBalance > 0.010) or (RemainingBalance < -0.010)): RemainingBalance = InitialBalance for i in range(12): RemainingBalance = (RemainingBalance*(1 + MonthlyRate)) - Payment if RemainingBalance > 0.010: PaymentLow = Payment Payment = (Payment + PaymentHigh)/2.0 if RemainingBalance < -0.010: PaymentHigh = Payment Payment = (Payment + PaymentLow)/2.0 print ('Monthly payment to pay off debt in 1 year: ' + str(round(Payment,2))) print ('Balance: ' + str(round(RemainingBalance,2)))

OpenStudy (turingtest):

your formula to update the balance is wrong: RemainingBalance = (RemainingBalance*(1 + MonthlyRate)) - Payment payment needs to be subtracted before multiplication by 1+MonthlyRate

OpenStudy (turingtest):

that link is not working for me, but I did the same problem once on OCW and once for edX so I think mine is right did you try it to see if you get the right answer

OpenStudy (turingtest):

?

OpenStudy (anonymous):

No it doesn't give the correct answer and it doesn't jive with interest payments. The interest accrues and then you pay off the interest plus a portion of the principle. I also checked my answer in excel and it is correct.

OpenStudy (turingtest):

what are the inputs for balance and interest that you are trying to get the answer 29643.05 from?

OpenStudy (anonymous):

Ah that's it, I was entering in .02 instead of .2 -- I did the program right. Thanks!

OpenStudy (turingtest):

congrads lol